a crow left Valley College at 2:00 pm and flew straight east at 21 miles per hour. another crow left Valley College at 3:oo pm and flew straight north at 24 miles per hour. how far apart (as the crow flies) are they at 6:00 pm (to the nearest tenth of a mile)
Answer by TimothyLamb(4379)

s = d / t
d = s * t
---
east crow at 6 pm:
t = 6 - 2 = 4 hours
d = 21 * 4 = 84 miles
---
north crow at 6 pm:
t = 6 - 3 = 3 hours
d = 24 * 3 = 72 miles
---
distance between crows at 6 pm:
use pythagorean thm:
h = sqrt( aa + bb )
h = sqrt( 84^2 + 72^2 )
h = sqrt( 7056 + 5184 )
h = sqrt( 12240 )
---
h = 110.6 miles
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php